Below are the two most recent balance sheets and most recent income statement for Seashell Industries. The company has an effective tax rate of 35% Balance Sheet 2018 2017 Assets Cash Accounts Receivable (net) Inventory Long-lived assets Less: Accumulated depreciation $10,000 6,000 1,500 10,000 11,000 (4.000 2,000) S32.000 S26.500 6,000 8,000 12,000 Total assets Liabilities and Stockholders Equity: Accounts payable Deferred revenues Long-term note payable Less: Discount on note payable Common stock Retained earnings S 5,000 1,000 10,000 (800) 12,000 4,800 S 6,000 2,000 10,000 (1,000) 6,000 3,500 S32.000 S26500 Total liabilities and stockholders equity Income Statement For the year ended December 31, 2018 Revenues Cost of goods sold Depreciation expense Interest expense Bad debt expense Other expense (including income taxes) Net income $42,000 (24,000) (2,000) (3,000) (2,000) (9.000) S 2,000 Calculate the following financial ratios for 2018 (1) Accounts receivable turnover
Answer #1

Ans. Accounts receivable turnover ratio = Revenue / Average accounts receivable

= 42000 / 3750

= 11.20 times

* Average accounts receivable = (beginning accounts receivable + ending accounts receivable) / 2

= (1500 + 6000) / 2

= 7500 / 2

= 3750